# Performance of Practical Quantum Oblivious Key Distribution

### v2

@@ VI Discussion
Classical OT protocols based on asymmetric cryptography comprise the overwhelming majority of current real-world implementations of OT. The obvious main advantage of quantum OT is the weaker computational hardness assumption (OWF vs asymmetric cryptography), while the main advantage of current post-quantum classical OT implementations is speed. As shown in Fig. 4, the presented experimental setup is able to produce up to 0.100.10~{}0.100.10 OT/s, which pales in comparison to contemporary classical protocols, such as [32, 33, 34, 35], that can achieve upwards of $10^{5}$ OT/s (not including latency between parties) with current off-the-shelf hardware (for more details, see [35]). This difference can be mitigated by the use of OT extension algorithms, as the difference in speed would only matter during the generation of the base OTs. Note that in this case one should use a OT extension that matches the computational assumption of this work, such as [49].
@@
Quantum protocols, both discrete variable (DV) [29] and continuous variable (CV) [31], have been shown to achieve statistically-secure OT in the noisy-storage model. Their experimental implementations show comparable values of quantum communication cost in terms of shared signals: $10^{8}$ (no memory encoding assumption), and $10^{5}$ (Gaussian encoding) for CV, and $10^{7}$ for DV. As shown in Fig. 3, our protocol requires $10^{6}$ quantum signals when matching their security ($\varepsilon=10^{-7}$), which improves upon the alternatives when no additional assumption on the memory encoding of the adversary is made. Less straightforward to compare is the strength of the assumptions of noisy storage and OWFs. We note that the existence of OWFs is an assumption that permeates modern cryptography, from block cipher encryption and message authentication up to public-key cryptography protocols [50], which makes $\pi_{\textnormal{QROT}}$ more suited to be introduced in current cipher suites than protocols with alternative assumptions. In particular, as noted above, OWFs are required for OT extension algorithms.
Regarding potential improvements and further work, we can identify two main directions to build upon this work: performance and security. Regarding performance, we note that dominant term in the expression for $\varepsilon_{max}$ is the one associated with the significance of the parameter estimation (the first term in Eq. 29). This translates into the relatively large values of $N_{0}$ needed to achieve adequate security, which was the bottleneck in the performance of our implementation. One way to reduce the number of signals needed per OT is to modify the protocol to perform many concurrent ROTs in a single run. This would mean performing one single estimation, albeit of a larger sample, that would work for many OTs in such a way that the required number of signals per ROT is decreased. On the topic of increasing security it seems like a natural step to transform our indistinguishability-based security definition into a simulation-based one. This would involve replacing the commitment scheme with one that holds the properties of equivocability and extractability. Such constructions based only on OWF have been developed in [40] and independently in [51]. Although using these type of commitment schemes can non-trivially increase the complexity of the protocol, the composability properties associated with them are considered desirable in the context of using the OTs as a resource for arbitrary MPC.
VII Acknowledgements
@@
Appendix A Detailed proof of Theorem III.1
A.1 Supporting lemmas
@@
$\begin{split}||\rho_{T,\hat{X},X,E}-\tilde{\rho}_{T,\hat{X},X,E}||^{2}&\leq\sum_{I_{1},I_{2}}q(I_{1},I_{2}){\lambda^{\perp}_{I_{1},I_{2}}}^{2}\\ &=\sum_{I_{1},I_{2}}q(I_{1},I_{2})\sum_{x\notin B_{I_{1},I_{2}}}|\beta_{x}|^{2}\\ &=\sum_{I_{1},I_{2}}q(I_{1},I_{2})\sum_{x\in\{0,1\}^{N}}\xi(I_{1},I_{2},x)|\beta_{x}|^{2}\\ &=\sum_{x\in\{0,1\}^{N}}|\beta_{x}|^{2}\sum_{I_{1},I_{2}}q(I_{1},I_{2})\xi(I,x)\\ &\leq\sum_{x\in\{0,1\}^{N}}|\beta_{x}|^{2}\mathcal{Q}(N,\varepsilon)=\mathcal{Q}(N,\varepsilon),\end{split}$ (64)
@@
A.2 Proof of Lemma IV.1
Here we present a proof of Lemma IV.1 used in the protocol’s correctness analysis in Section IV.
Lemma A.4.
@@
Now that we have a form for the conditioned state as pointed out in Eq. (A.2), we can move to the action of Step (11), where Bob computes $y^{B}=\mathtt{dec}(\mathtt{syn}(x^{A}_{I_{0}},x^{B}_{I_{0}}))$. The resulting state of the systems $C,X^{A}_{I_{0}},X^{A}_{I_{1}},Y^{B}$ is then given by:
@@
[…]
@@
A.3 Proof of Lemma IV.2
Here we present a proof of the Lemma IV.2 introduced in Section IV.2 regarding the hiding property of the commitment in the context of $\pi_{\textnormal{QROT}}$.
Lemma A.5.
@@
After uniformly sampling the values of $\theta^{B}$, Bob proceeds to perform his measurement on his qubit shares. Let $\mathcal{H}_{X^{B}}$ denote the system where Bob records the outcome string $x^{B}$. Additionally, at Step (3) Bob receives the value of $r$, this is a classical message, which we model as Bob receiving the $\mathcal{H}_{R}$ system and measuring it in the computational basis upon arrival. We can now easily use Eq. (A.3) to get the post-measurement state at the end of Step (3) after tracing out $\mathcal{H}_{\Phi^{B}}$, which is given by
@@
Let $J\subseteq I$, we now want to use the hiding property of the commitment scheme to approximate the state (A.3) to one where the values of com and $\textnormal{open}_{J}$ don’t provide any information about $\theta^{B}_{\bar{J}}$. First, we proceed to rewrite the expression for the $\textnormal{COM}_{i}$ and $\textnormal{OPEN}_{i}$ subsystems by grouping the individual values of $\textnormal{com}_{i}$
@@
where $P^{\theta_{i},x_{i},r}_{\textnormal{com}}$ is the respective distribution for $\textnormal{com}_{i}$ for uniformly sampled $s_{i}$, which depends on the commitment scheme used, and $\mathcal{C}_{r}(\textnormal{com}_{i})$ is the set of strings $\textnormal{open}_{i}$ that satisfy $\mathtt{ver}(\textnormal{com}_{i},\textnormal{open}_{i},r)\neq\bot$. Substituting Eq. (A.3) into Eq. (A.3) and tracing out $\textnormal{OPEN}_{\bar{J}}$ and $R$ results in
@@
Applying Eq. (106) to the $\bar{J}$ subset in Eq. (105), and from Lemma II.2 (2) and (3) we get that
@@
Finally, using Lemma II.2 (1) and (2) we obtain the required result
@@
A.4 Proof of Lemma IV.3
In this section we present a proof of Lemma IV.3, which states that the string separation step of $\pi_{\textnormal{QROT}}$ does not leak any information about the random bit $c$ to the receiver.
Lemma A.6.
@@
where the sum in SEP goes over all possible $I_{0},I_{1}$ given $I_{t}$. Tracing out $\Theta^{B}_{\bar{I}_{t}}$ and using Eq. (A.4) we obtain
@@
A.5 Proof of Lemma IV.4
In this section we present a proof of Lemma IV.4, introduced in Section IV as part of the security analysis against a dishonest receiver. Recall that the transcript of the protocol $\vec{\tau}=(x^{A}_{I_{t}},\theta^{A},r,\textnormal{com},I_{t},I_{s},\textnormal{open}_{I_{t}},r)$ is defined to consist of all classical information (with the exception of her measurement outcomes) that Alice has access up to Step (8) of the protocol.
Lemma A.7.
@@
We proceed by tracking the properties of Alice’s and Bob’s shared state as the protocol develops in order to bound the conditional min-entropy of Alice’s measurement outcomes given the information the Bob gains during the protocol, then we use Lemma II.4 to obtain the desired result. Let $\rho^{\textnormal{rand}}_{\Theta^{A},T,R}$ denote the quantum state associated to the systems holding Alice’s basis choice $\theta^{A}$, test subset $I_{t}$, and the value of $r$ used in the commit/open phase, which we can treat as if they are sampled at the beginning of the protocol since their distribution is fixed, and is given by
@@
We intend to use Lemma A.1 to bound the form of the shared state after the parameter estimation step, and then Lemma A.3 to bound the amount of correlation between Alice’s measurement outcomes on the system $\Phi^{A}$ and Bob’s system. For that, we first need to associate Bob’s commitments with their corresponding committed strings $x^{B}$ and $\theta^{B}$. For an arbitrary dishonest Bob the strings that Alice received are not guaranteed to be outcomes of the $\mathtt{com}$ function and may not have an associated preimage. Consider now the functions $x^{B}_{i}(r,\textnormal{com}),\theta^{B}_{i}(r,\textnormal{com}):\{0,1\}^{n_{r}}\times\{0,1\}^{n_{c}}\rightarrow\{0,1\}$ defined as follows,
@@
$\theta^{B}_{i}(r,\textnormal{com})=\begin{cases}\theta&\quad\textnormal{if }\textnormal{com}_{i}=\mathtt{com}((\theta,x),s,r)\textnormal{ for some }x\in\{0,1\},s\in\{0,1\}^{n_{s}}\\ 0&\quad\textnormal{otherwise }\end{cases},$ (130)
@@
Recall that $I_{s}(\theta^{A},\theta^{B},I_{t})=\{i\in I_{t}:\theta^{A}_{i}\oplus\theta^{B}_{i}=0\}$. From Lemma A.1 we know that there exists a state
@@
×$\times\gamma^{x,\theta^{B},r,\textnormal{com},I_{s},I_{t}}{|{\tilde{\phi}^{x,\theta^{A},r,\textnormal{com},I_{t}}}\rangle}_{B_{2}}{|{x_{\bar{I}_{t}},\theta^{B}_{\bar{I}_{t}}}\rangle}_{\Phi^{A}_{\bar{I}_{t}}},$ (143)
@@
During Step (8), Alice sends the $\Theta^{A}$ system to Bob, who then computes $J_{0},J_{1}$ (in the actual protocol, Alice sends only $\Theta^{A}_{{\bar{I_{t}}}}$, but to simplify the expressions we can assume, without loss of generality, that she sends the whole register $\Theta^{A}$). To simplify the list of dependencies, denote the transcript of the protocol up until Step (8) as $\vec{\tau}=(x^{A}_{I_{t}},\theta^{A},r,\textnormal{com},I_{t},I_{s},\textnormal{open}_{I_{t}})$. Keep in mind that, although $\vec{\tau}$ consists of seven quantities, $I_{s}$ and $\textnormal{open}_{I_{t}}$ are completely defined by the other five. In the remaining of the proof, unless noted otherwise, the sums over $\vec{\tau}$ run over the values of its variables as shown in Eq. (A.5). By defining
@@
Additionally, for any given $J_{0},J_{1}$, denote by $J_{d}$ the complement of $J_{0}\cup J_{1}$ in $\bar{I}_{t}$. Following Lemma II.3 (3) and (5) we know that for any $b\in\{0,1\}$
@@
We can invoke Lemma A.3 to obtain an expression for the above quantity explicitly in terms of the protocol parameters $N_{0},\alpha,\delta_{1}$, and $\delta_{2}$. For that, we must take a small detour to define the associated mixed states $\rho^{\textnormal{mix}}_{\Phi^{A}_{{J_{b}}}B_{3}}$ and compute their respective post-measurement entropy. First, for $b\in\{0,1\}$, we compute the reduced states
@@
$B_{b}(x_{{J_{\bar{b},d}}})$ =$=\{x_{{J_{b}}}:x_{{J_{b,\bar{b},d}}}\in G(x^{A}_{I_{s}},r,\textnormal{com})\}$
=$=\{x_{{J_{b}}}:|(\frac{1}{2}-\delta_{2})r_{\textnormal{H}}(x_{{J_{b}}}\oplus x^{B}_{{J_{b}}})+(\frac{1}{2}+\delta_{2})r_{\textnormal{H}}(x_{{J_{\bar{b},d}}}\oplus x^{B}_{{J_{\bar{b},d}}})-r_{\textnormal{H}}(x^{A}_{I_{s}}\oplus x^{B}_{I_{s}})|\leq\delta_{1}\},$ (159)
@@
which is a product state between the systems $X^{A}_{{J_{b}^{1}}}$ and $X^{A}_{{J_{b}^{0}}}B_{3}$. From Lemma II.3 (1) and (2) we know that
@@
Application of Lemma A.3 together with equations (A.5) and (160) leads to
@@
Note that the above expression depends only on the number of nonmatching bases $\theta^{\textnormal{ch}}$ associated to the indices in $J_{b}$ and the parameters of the protocol, which in turn makes the infimum in Eq. (A.5) straightforward to compute. We can now add the respective conditional min-entropies for $X^{A}_{J_{0}}$ and $X^{A}_{J_{1}}$, which results in:

### v3

@@ 6 Discussion
Classical OT protocols based on asymmetric cryptography comprise the overwhelming majority of current real-world implementations of OT. The obvious main advantage of quantum OT is the weaker computational hardness assumption (OWF vs asymmetric cryptography), while the main advantage of current post-quantum classical OT implementations is speed. As shown in Fig. 5, the presented experimental setup is able to produce up to 0.10~0.10 OT/s, which pales in comparison to contemporary classical protocols, such as [30, 31, 32, 33], that can achieve upwards of $10^{5}$ OT/s (not including latency between parties) with current off-the-shelf hardware (for more details, see [33]). This difference can be mitigated by the use of OT extension algorithms, as the difference in speed would only matter during the generation of the base OTs. Note that in this case one should use a OT extension that matches the computational assumption of this work, such as [54].
@@
Quantum protocols, both discrete variable (DV) [27] and continuous variable (CV) [29], have been shown to achieve statistically-secure OT in the Quantum Noisy-Storage model (QNS). Their experimental implementations show comparable values of quantum communication cost in terms of shared signals: $10^{8}$ (no memory encoding assumption), and $10^{5}$ (Gaussian encoding) for CV, and $10^{7}$ for DV. As shown in Fig. 4, our protocol requires $10^{6}$ quantum signals when matching their security ($\varepsilon=10^{-7}$), which improves upon the alternatives when no additional assumption on the memory encoding of the adversary is made. Less straightforward to compare is the strength of the assumptions of noisy storage and OWFs. We note that the existence of OWFs is an assumption that permeates modern cryptography, from block cipher encryption and message authentication up to public-key cryptography protocols [55], which makes $\pi_{\textnormal{QROT}}$ more suited to be introduced in current cipher suites than protocols with alternative assumptions. In particular, as noted above, OWFs are required for OT extension algorithms. A summary of comparisons between the different approaches can be found in Fig.2.
Protocol Type Assumption Quantum Cost Security
This work Quantum Discrete Variable OWF $\mathcal{O}(N)$ Indistinguishability UC ROM
GLSV21 [38]* Quantum Discrete Variable OWF $Poly(N)$ Stand-alone plain model
S10 [47, 27] Quantum Discrete Variable QNS $\mathcal{O}(N)$ Indistinguishability
FGSPSW18 [29] Quantum Continuous Variable QNS $\mathcal{O}(N)$ Indistinguishability
MR19 [30] Classical DDH Stand-alone ROM
BFGMMS21 [33] Classical RLWE UC ROM
P16 [56]* Quantum/Relativistic Discrete Variable SLS $\mathcal{O}(N)$ Other
Table 2: Comparison of our work with other approaches for OT. NN denotes the respective security parameter. Acronyms for assumptions are as follows: OWF - One Way Functions; QNS - Quantum Noisy Storage; DDH - Decisional Diffie-Helmann; RLWE - Ring Learning With Errors - SLS - Space-Like Separation enforced. Protocols marked with * do not have a reference experimental implementation at the time of writing.
Regarding potential improvements and further work, we can identify two main directions to build upon this work: performance and security. Regarding performance, we note that dominant term in the expression for $\varepsilon_{max}$ is the one associated with the significance of the parameter estimation (the first term in Eq. 13). This translates into the relatively large values of $N_{0}$ needed to achieve adequate security, which was the bottleneck in the performance of our implementation. One way to reduce the number of signals needed per OT is to modify the protocol to perform many concurrent ROTs in a single run. This would mean performing one single estimation, albeit of a larger sample, that would work for many OTs in such a way that the required number of signals per ROT is decreased. On the topic of increasing security two main directions come to mind. First, we can consider the constructions of collapsing hash functions proposed in [57, 58] to implement statistically hiding, computationally collapse binding commitments, which in turn allow for OT protocols that feature forward security (the OT remains secure even if the underlying hash function can be attacked after the commit/open phase of the protocol). The second direction would be a deeper exploration of the composable security of the protocol in the ROM. This can come from generalizing Theorem 4.1 for any weakly-interactive commitments (currently the proof applies only to the LRV25 construction), or applying the techniques developed in [59] to prove UC security of commitments in the quantum ROM to remove the adversary’s limitation of classical access to the oracle. From the practical implementation perspective, it seems natural to integrate quantum OT into both QKD setups for a unified physical layer capable of providing secure communication and computation powered by OT extension and MPC algorithms, bringing the benefits of quantum OT closer to real world usage.
7 Acknowledgements
@@
Appendix A Preliminaries
A.1 Quantum computational efficiency and distinguishability
We model the quantum capabilities of parties through programs running on quantum computers, for which we adopt a model based on deterministic-control quantum Turing Machines [60]. For the purposes of the following definitions, a quantum computer is a device that has a classical interface and a quantum part, which contains the quantum memory registers available to the party. The classical interface has the capabilities of a classical computer augmented with the ability to perform a predefined universal set of quantum operators on the quantum memory registers and perform measurements in the canonical (computational) basis. Given a specified type of quantum computer, a quantum program is a classical description of a set of instructions to be run by the computer, including the quantum operations and measurements to be executed in the quantum part, as well as any classical computation. Quantum programs can be compared with probabilistic classical programs as they both have natural numbers as inputs/outputs. When a quantum computer runs the program $T$ with input $x\in\mathbb{N}$, we assume that the quantum part of the computer starts with some predefined initial state, performs a sequence of operations on its quantum registers, and upon halting, it outputs $T(x)\in\mathbb{N}$ on its classical interface by reading the appropriate registers associated with the program’s output. Each execution of a quantum program is then associated to a quantum operation, which is the result of all the operations performed on the quantum part during the execution of the program.
Definition A.1.
(Computational efficiency)
Definition A.2.
(Distinguishing Advantage)
Definition A.3.
(Indistinguishability - Finite)
Proposition A.1.
$\rho_{1}\approx_{D(\rho_{1},\rho_{2})}\rho_{2}.$ (42)
Definition A.4.
(Indistinguishability – Asymptotical)
Recall from Section 2 that, when the parameter $k$ is implicit, we may omit the explicit dependence on $k$ and use $\approx$ and $\approx^{(c)}$ for statistical and computational indistinguishability, respectively. We now turn our attention to the properties of indistinguishable states. It is worth noting that computational indistinguishability is only meaningful in terms of information security when the adversary is assumed to have limited computational capabilities. It is important then to define the type of quantum operations such adversary can perform:
Definition A.5.
(Efficient quantum operation)
The following properties are straightforward to prove from Definitions A.3 and A.4 and the properties of trace distance:
Lemma A.1.
(Properties of indistinguishable states I)
Let $\rho_{1},\rho_{2},\rho_{3}\in\mathcal{D}(\mathcal{H})$:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Lemma A.2.
(Properties of indistinguishable states II)
  1. 1.
  2. 2.
  3. 3.
  4. 4.
A.2 Entropic quantities
Definition A.6.
(Max-entropy)
Definition A.7.
(Min-entropy and conditional min-entropy)
$H_{\textnormal{min}}(\rho)=-\log(\lambda_{\max}(\rho)).$ (46)
Definition A.8.
(Partially classical states)
Lemma A.3.
(Properties of min- and max-entropy)
Let $\varepsilon,\varepsilon^{\prime}\geq 0$:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
We use universal hashing to implement randomness extraction in the final steps of the protocol. The proof both Lemmas A.3 and A.4 can be found in [63].
Definition A.9.
(Universal hashing)
$\textnormal{Pr}[f_{i}(x)=f_{i}(y)]\leq\frac{1}{2^{n}}.$ (51)
Lemma A.4.
(Quantum leftover hash)
$\otimes|x\rangle\ \!\!\!\langle x|_{A}\rho^{x}_{B}\Big{)},$ (52)
it holds that
with
Appendix B Detailed proof of Theorem 3.1
B.1 Supporting lemmas
@@
[…]
@@
B.2 Proof of Lemma 4.1
Here we present a proof of Lemma 4.1 used in the protocol’s correctness analysis in Section 4.
Lemma B.4.
@@
Now that we have a form for the conditioned state as pointed out in Eq. (B.2), we can move to the action of Step (11), where Bob computes $y^{B}=\mathtt{dec}(\mathtt{syn}(x^{A}_{I_{0}},x^{B}_{I_{0}}))$. The resulting state of the systems $C,X^{A}_{I_{0}},X^{A}_{I_{1}},Y^{B}$ is then given by:
@@
for some coefficients $P_{\textnormal{correct}},P_{\bot},P_{\textnormal{error}}$, and state $\sigma$ orthogonal to both $|x^{A}_{I_{0}}\rangle\ \!\!\!\langle x^{A}_{I_{0}}|$ and $|\bot\rangle\ \!\!\!\langle\bot|$. By applying the verifiability property of the IR scheme with security parameter $k^{\prime}$ (where $P_{\textnormal{error}}=\varepsilon_{\textnormal{IR}}(k^{\prime})$), and conditioning the resulting state to not having aborted, we get the desired result
@@
B.3 Proof of Lemma 4.2
Here we present a proof of the Lemma 4.2 introduced in Section 4.2 regarding the hiding property of the commitment in the context of $\pi_{\textnormal{QROT}}$.
Lemma B.5.
@@
After uniformly sampling the values of $\theta^{B}$, Bob proceeds to perform his measurement on his qubit shares. Let $\mathcal{H}_{X^{B}}$ denote the system where Bob records the outcome string $x^{B}$. Additionally, at Step (3) Bob receives the value of $r$, this is a classical message, which we model as Bob receiving the $\mathcal{H}_{R}$ system and measuring it in the computational basis upon arrival. We can now easily use Eq. (B.3) to get the post-measurement state at the end of Step (3) after tracing out $\mathcal{H}_{\Phi^{B}}$, which is given by
@@
Let $J\subseteq I$, we now want to use the hiding property of the commitment scheme to approximate the state (B.3) to one where the values of com and $\textnormal{open}_{J}$ don’t provide any information about $\theta^{B}_{\bar{J}}$. First, we proceed to rewrite the expression for the $\textnormal{COM}_{i}$ and $\textnormal{OPEN}_{i}$ subsystems by grouping the individual values of $\textnormal{com}_{i}$
@@
where $P^{\theta_{i},x_{i},r}_{\textnormal{com}}$ is the respective distribution for $\textnormal{com}_{i}$ for uniformly sampled $s_{i}$, which depends on the commitment scheme used, and $\mathcal{C}_{r}(\textnormal{com}_{i})$ is the set of strings $\textnormal{open}_{i}$ that satisfy $\mathtt{ver}(\textnormal{com}_{i},\textnormal{open}_{i},r)\neq\bot$. Substituting Eq. (B.3) into Eq. (B.3) and tracing out $\textnormal{OPEN}_{\bar{J}}$ and $R$ results in
@@
Applying Eq. (106) to the $\bar{J}$ subset in Eq. (105), and from Lemma A.2 (2) and (3) we get that
@@
Finally, using Lemma A.2 (1) and (2) we obtain the required result
@@
B.4 Proof of Lemma 4.3
In this section we present a proof of Lemma 4.3, which states that the string separation step of $\pi_{\textnormal{QROT}}$ does not leak any information about the random bit $c$ to the receiver.
Lemma B.6.
@@
where the sum in SEP goes over all possible $I_{0},I_{1}$ given $I_{t}$. Tracing out $\Theta^{B}_{\bar{I}_{t}}$ and using Eq. (B.4) we obtain
@@
B.5 Proof of Lemma 4.4
In this section we present a proof of Lemma 4.4, introduced in Section 4 as part of the security analysis against a dishonest receiver. Recall that the transcript of the protocol $\vec{\tau}=(x^{A}_{I_{t}},\theta^{A},r,\textnormal{com},I_{t},I_{s},\textnormal{open}_{I_{t}},r)$ is defined to consist of all classical information (with the exception of her measurement outcomes) that Alice has access up to Step (8) of the protocol.
Lemma B.7.
@@
We proceed by tracking the properties of Alice’s and Bob’s shared state as the protocol develops in order to bound the conditional min-entropy of Alice’s measurement outcomes given the information the Bob gains during the protocol, then we use Lemma A.4 to obtain the desired result. Let $\rho^{\textnormal{rand}}_{\Theta^{A},T,R}$ denote the quantum state associated to the systems holding Alice’s basis choice $\theta^{A}$, test subset $I_{t}$, and the value of $r$ used in the commit/open phase, which we can treat as if they are sampled at the beginning of the protocol since their distribution is fixed, and is given by
@@
We intend to use Lemma B.1 to bound the form of the shared state after the parameter estimation step, and then Lemma B.3 to bound the amount of correlation between Alice’s measurement outcomes on the system $\Phi^{A}$ and Bob’s system. For that, we first need to associate Bob’s commitments with their corresponding committed strings $x^{B}$ and $\theta^{B}$. For an arbitrary dishonest Bob the strings that Alice received are not guaranteed to be outcomes of the $\mathtt{com}$ function and may not have an associated preimage. Consider now the functions $x^{B}_{i}(r,\textnormal{com}),\theta^{B}_{i}(r,\textnormal{com}):\{0,1\}^{n_{r}}\times\{0,1\}^{n_{c}}\rightarrow\{0,1\}$ defined as follows,
@@
[…]
@@
Recall that $I_{s}(\theta^{A},\theta^{B},I_{t})=\{i\in I_{t}:\theta^{A}_{i}\oplus\theta^{B}_{i}=0\}$. From Lemma B.1 we know that there exists a state
@@
$\times\gamma^{x,\theta^{B},r,\textnormal{com},I_{s},I_{t}}{|{\tilde{\phi}^{x,\theta^{A},r,\textnormal{com},I_{t}}}\rangle}_{B_{2}}{|{x_{\bar{I}_{t}},\theta^{B}_{\bar{I}_{t}}}\rangle}_{\Phi^{A}_{\bar{I}_{t}}},$ (143)
@@
During Step (8), Alice sends the $\Theta^{A}$ system to Bob, who then computes $J_{0},J_{1}$ (in the actual protocol, Alice sends only $\Theta^{A}_{{\bar{I_{t}}}}$, but to simplify the expressions we can assume, without loss of generality, that she sends the whole register $\Theta^{A}$). To simplify the list of dependencies, denote the transcript of the protocol up until Step (8) as $\vec{\tau}=(x^{A}_{I_{t}},\theta^{A},r,\textnormal{com},I_{t},I_{s},\textnormal{open}_{I_{t}})$. Keep in mind that, although $\vec{\tau}$ consists of seven quantities, $I_{s}$ and $\textnormal{open}_{I_{t}}$ are completely defined by the other five. In the remaining of the proof, unless noted otherwise, the sums over $\vec{\tau}$ run over the values of its variables as shown in Eq. (B.5). By defining
@@
Additionally, for any given $J_{0},J_{1}$, denote by $J_{d}$ the complement of $J_{0}\cup J_{1}$ in $\bar{I}_{t}$. Following Lemma A.3 (3) and (5) we know that for any $b\in\{0,1\}$
@@
We can invoke Lemma B.3 to obtain an expression for the above quantity explicitly in terms of the protocol parameters $N_{0},\alpha,\delta_{1}$, and $\delta_{2}$. For that, we must take a small detour to define the associated mixed states $\rho^{\textnormal{mix}}_{\Phi^{A}_{{J_{b}}}B_{3}}$ and compute their respective post-measurement entropy. First, for $b\in\{0,1\}$, we compute the reduced states
@@
$B_{b}(x_{{J_{\bar{b},d}}})$ $=\{x_{{J_{b}}}:x_{{J_{b,\bar{b},d}}}\in G(x^{A}_{I_{s}},r,\textnormal{com})\}$
$=\{x_{{J_{b}}}:|(\frac{1}{2}-\delta_{2})r_{\textnormal{H}}(x_{{J_{b}}}\oplus x^{B}_{{J_{b}}})+(\frac{1}{2}+\delta_{2})r_{\textnormal{H}}(x_{{J_{\bar{b},d}}}\oplus x^{B}_{{J_{\bar{b},d}}})-r_{\textnormal{H}}(x^{A}_{I_{s}}\oplus x^{B}_{I_{s}})|\leq\delta_{1}\},$ (159)
@@
which is a product state between the systems $X^{A}_{{J_{b}^{1}}}$ and $X^{A}_{{J_{b}^{0}}}B_{3}$. From Lemma A.3 (1) and (2) we know that
@@
Application of Lemma B.3 together with equations (B.5) and (160) leads to
@@
Note that the above expression depends only on the number of nonmatching bases $\theta^{\textnormal{ch}}$ associated to the indices in $J_{b}$ and the parameters of the protocol, which in turn makes the infimum in Eq. (B.5) straightforward to compute. We can now add the respective conditional min-entropies for $X^{A}_{J_{0}}$ and $X^{A}_{J_{1}}$, which results in: